first order rate constant for decomposition of N2O4 is 6.2*10^-4 per sec. half period for this decomposition is

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
for a first order reaction
 
t(1/2) = ln2 / k
 
t(1/2) = 0.693 / 6.2 * 10^-4
 
 = 1117 sec
 
Regards
Arun (askIITians forum expert)